# Can Watermarking Large Language Models Prevent Copyrighted Text Generation and Hide Training Data?

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in generating diverse and contextually rich text. However, concerns regarding copyright infringement arise as LLMs may inadvertently produce copyrighted material. In this paper, we first investigate the effectiveness of watermarking LLMs as a deterrent against the generation of copyrighted texts. Through theoretical analysis and empirical evaluation, we demonstrate that incorporating watermarks into LLMs significantly reduces the likelihood of generating copyrighted content, thereby addressing a critical concern in the deployment of LLMs. Additionally, we explore the impact of watermarking on Membership Inference Attacks (MIAs), which aim to discern whether a sample was part of the pretraining dataset and may be used to detect copyright violations. Surprisingly, we find that watermarking adversely affects the success rate of MIAs, complicating the task of detecting copyrighted text in the pretraining dataset. Finally, we propose an adaptive technique to improve the success rate of a recent MIA under watermarking. Our findings underscore the importance of developing adaptive methods to study critical problems in LLMs with potential legal implications.
1 Introduction
Preprint. Correspondence to: Michael-Andrei Panaitescu-Liess ¡mpanaite@umd.edu¿.
@@ 1 Introduction
Current lawsuits against AI companies for unauthorized use of copyrighted content (e.g., Andersen v. Stability AI Ltd, NYT v. OpenAI) highlight the urgent need for methods to address these challenges. In this paper, we focus on watermarking LLMs to tackle two main issues: (1) preventing the generation of copyrighted content, and (2) detecting copyrighted content in training data. We show that watermarking can significantly impact both the generation of copyrighted text and the detection of copyrighted content in training data.
Firstly, we observe that current LLM output watermarking techniques can significantly reduce the probability of LLMs generating copyrighted content, by tens of orders of magnitude. Our empirical results focus on two recent watermarking methods: UMD [20] and Unigram-Watermark [43]. Both methods split the vocabulary into two sets (green and red) and bias the model towards selecting tokens from the green set by altering the logits distribution, thereby embedding a detectable signal. We provide both empirical and theoretical results to support our findings.
@@
Finally, we propose an adaptive method designed to enhance the success rate of a recent MIA [31] in detecting copyright violations under watermarking. This method applies a correction to the model’s output to account for the perturbations introduced by watermarks. By incorporating knowledge about the watermarking scheme, we improve the detection performance for pretraining data, counteracting the obfuscation caused by watermarking. Our contribution underscores the importance of continuously developing adaptive attack methodologies to keep pace with advances in defense mechanisms.
@@
The rest of the paper is organized as follows. In Section 2, we discuss prior work on LLM watermarking, copyright, memorization, and membership inference. We formally introduce the problems that we study in Section 3 and present the first two contributions of empirical results and theoretical analysis in Sections 4 and 5. In Section 6, we introduce the adaptive version of the Min-KProb membership inference attack. Finally, in Section 7, we discuss the limitations of our work and provide concluding remarks. Additional experiments as well as proofs are included in Appendix A.
Figure 1: Illustration of the effect of LLM watermarking on generation of copyrighted content. We observe that watermarking can make it more than $10^{20}$ times less likely for Llama-30B to generate copyrighted content.
2 Related work
@@
Watermarks for LLMs. Language model watermarking techniques embed identifiable markers into output text to detect AI-generated content. Recent strategies incorporate watermarks during the decoding phase of language models [43, 20]. Aaronson [1] develops the Gumbel watermark, which employs traceable pseudo-random sampling for generating subsequent tokens. Kirchenbauer et al. [20] splits the vocabulary into red and green lists according to preceding tokens, biasing the generation towards green tokens. Zhao et al. [43] employs a fixed grouping strategy to develop a robust watermark with theoretical guarantees. Liu et al. [24] proposes to generate watermark logits based on the preceding tokens’ semantics rather than their token IDs to boost the robustness of the watermark. Kuditipudi et al. [22] and Christ et al. [10] explore watermark methods that do not change the output textual distribution.
@@
Copyright. Copyright protection in the age of AI has gained importance, as discussed by Ren et al. [28]. Vyas et al. [38] addresses content protection through near access-freeness (NAF) and developed learning algorithms for generative models to ensure compliance under NAF conditions. Prior works focus on training algorithms to prevent copyrighted text generation [38, 11], whereas our work emphasizes lightweight, inference-time algorithms. Other works have studied copyright in machine learning from a legal perspective. Hacohen et al. [15] utilizes a generative model to determine the generic characteristics of works to aid in defining the scope of copyright. Elkin-Koren et al. [14] demonstrates that copying does not necessarily constitute copyright infringement and argues that existing detection methods may detract from the foundational purposes of copyright law.
Memorization. One cause of copyright issues is that machine learning models may memorize training data. Prior studies have observed that LLMs can memorize copyrighted or private information in training data, such as phone numbers and addresses [19, 5, 6, 23], leading to significant privacy and security concerns. To measure memorization, Carlini et al. [6] proposes eidetic memorization, defining a string as memorized if it was present in the training data and it can be reproduced by a prompt. This definition, along with variations like exact and perfect memorization, has been widely adopted in subsequent studies [36, 18]. Carlini et al. [8] quantitatively measures memorization in LLMs as the fraction of extractable training data and finds that memorization significantly grows as model size scales and training examples are duplicated. To minimize memorization, Lee et al. [23] and Kandpal et al. [18] propose deduplicating training data, which also improves accuracy. Hans et al. [16] proposes the Goldfish Loss as a training-time defense against verbatim memorization. Ippolito et al. [17] proposes an inference time defense that perfectly prevents all verbatim memorization. However, it cannot prevent the leakage of training data due to the existence of many “style-transfer” prompts, suggesting it is a challenging open problem. Unlike the methods that we are studying in this paper, Ippolito et al. [17] requires access to a complete set of copyrighted texts that the model was trained on. Memorization in the image domain has also been studied from various angles [33, 34, 9, 39].
Membership Inference. As a proxy for measuring memorization, membership inference attacks (MIAs) predict whether or not a particular example was used to train the model [32, 40, 2]. Most membership inference attacks rely only on the model’s loss since the model is more likely to overfit an example if it is in the training data [29]. Carlini et al. [7] trains shadow models to predict whether an example is from the training data. In the NLP domain, many works have focused on masked language models [26] and fine-tuning data detection [35, 30]. Recently, Shi et al. [31] studies pretraining data inference and introduced a detection method based on the hypothesis that unseen examples are likely to contain outlier words with low probabilities under the LLM. Zhang et al. [41] approaches pretraining data detection by measuring how sharply peaked the likelihood is around the inputs. Duarte et al. [13] proposes detecting copyrighted content in training data by probing the LLM with multiple-choice questions, whose options include both verbatim text and their paraphrases. Other methods include testing perplexity differences [25] and providing provable guarantees of test set contamination without access to pretraining data or model weights [27].
3 Setup and Notations
3.1 Definitions
@@
Verbatim Memorization of Copyrighted Content. For a fixed $k\in\mathbb{N}$, Carlini et al. [8] defines a string $s$ as being memorized by a model if $s$ is extractable with a prompt $p$ of length $k$ using greedy decoding and the concatenation $p\oplus s\in D$. We adopt a similar definition for verbatim memorization of copyrighted content but employ a continuous metric to measure it. Specifically, we measure verbatim memorization of a text $c\in C$ using the perplexity of the model on the copyrighted text $c_{p}$ when given the prefix $p$ as a prompt (where $c_{p}$ represents the text $c$ after removing its prefix $p$). Note that for $c_{p}=c_{p}^{(1)}\oplus c_{p}^{(2)}\oplus\dots\oplus c_{p}^{(n)}$ we compute the perplexity using the following formula $\texttt{perplexity}(c_{p}|p)=\big{(}\prod_{i=1}^{n}\mathbb{P}(c_{p}^{(i)}|p\oplus c_{p}^{(0)}\oplus c_{p}^{(1)}\oplus\dots\oplus c_{p}^{(i-1)}\big{)}^{-\frac{1}{n}}$, where $c_{p}^{(0)}$ is the empty string. In our experiments, $p$ is either an empty string or the first $10$, $20$, or $100$ tokens of $c$. Lower perplexity values thereby indicate higher levels of verbatim memorization.
MIAs for Copyrighted Training Data Detection. MIAs are privacy attacks aiming to detect whether a sample was part of the training set. We define an MIA for copyrighted data as a binary classifier $A(\cdot)$, which ideally outputs $A(x)=1,\forall x\in C_{D}$ and $A(x)=0,\forall x\in C-C_{D}$. In practice, $A(\cdot)$ is defined by thresholding a metric (e.g., perplexity), i.e., $A(x)=1,\forall x$ such that perplexity$(x)<t$ and 00, otherwise. Since the threshold $t$ needs to be set, prior work [31] uses AUC (Area Under the ROC Curve) as an evaluation metric which is independent of $t$. Note that we employ the same metric in our experiments.
LLM Watermarking. Watermarking LLMs consists of introducing signals during its training or inference that are difficult to detect by humans without the knowledge of a watermark key but can be detected using an algorithm if the key is known. We focus our paper on recent methods that employ logits distribution changes as a way of inserting watermark signals during the decoding process [20, 43]
3.2 MIAs
@@
Smaller Ref [6]. It is defined as the ratio of the log-perplexity of the target LLM on a sample to the log-perplexity of a smaller reference LLM on the same sample.
Lowercase [6]. This heuristic represents the ratio of the log-perplexity of the target LLM on the original sample to the log-perplexity of the LLM on the lowercase version of the sample.
Zlib [6]. It is defined as the ratio of the log-perplexity of the target LLM on a sample to the zlib entropy of the same sample.
@@
Min-KProb [31]. This heuristic computes the average of the minimum $K\%$ token probabilities outputted by the LLM on the sample. Note that this method requires tuning $K$, so in all our experiments we chose the best result over $K\%\in\{5\%,10\%,20\%,30\%,40\%,50\%,60\%\}$.
3.3 LLM Watermarking Methods
@@
UMD [20] splits the vocabulary into two sets (green and red) and biases the model towards the green tokens by altering the logit distribution. The hash of the previous token’s ID serves as a seed for a pseudo-random number generator used to split the vocabulary into these two groups. For a “hard” watermark, the model is forced not to sample from the red list at all. For a “soft” watermark, a positive bias $\delta$ is added to the logits of the green tokens before sampling. We focus our empirical evaluation on “soft” watermarks as they are more suitable for LLM deployment due to their smaller impact on the quality of the generated text.
@@
Unigram-Watermark [43] employs a similar approach of splitting the vocabulary into two sets and biasing the model towards one of the two sets. However, the split remains consistent throughout the generation of tokens. This choice is made to provide a provable improvement against paraphrasing attacks [21].
4 Watermarking LLMs Prevents Copyrighted Text Generation
@@
In this section, we study the effect of watermarking techniques on verbatim memorization. We discuss the implications of watermarking for preventing copyright text generation.
@@
Datasets. We consider $4$ versions of the WikiMIA benchmark [31] with 32, 64, 128, and 256 words in each sample and only consider the samples that were very likely part of the training set of all the models we consider (labeled as $1$ in Shi et al. [31]). We consider these subsets as a proxy for text that was used in the training set, and the model may be prone to verbatim memorization. From now on, we refer to this subset as the “training samples” or “training texts”. Similarly, we consider BookMIA dataset [31], which contains samples from copyrighted books.
@@
Metric. We measure the relative increase in perplexity on the generation of training samples by the watermarked model compared to the original model. We report the increase in both the minimum and average perplexity over the training samples. Note that a large increase in perplexity corresponds to a large decrease in the probability of generating that specific sample, as shown later in this section. When computing the perplexity, we prompt the model with an empty string, the first 10, and the first 20 tokens of the targeted training sample, respectively. In the BookMIA dataset, we designate the initial 100 tokens as the prompt. This is because each BookMIA sample contains 512 words, which is larger than the sample size in WikiMIA.
@@
Models. We conduct our empirical evaluation on $5$ recent LLMs: Llama-30B [37], GPT-NeoX-20B [4], Llama-13B [37], Pythia-2.8B [3] and OPT-2.7B [42].
4.1 Empirical Evaluation
Llama-30B NeoX-20B Llama-13B Pythia-2.8B OPT-2.7B
P. Min. Avg. Min. Avg. Min. Avg. Min. Avg. Min. Avg.
0 $3.3$ $31.2$ $3.7$ $52.1$ $4.9$ $34.3$ $11.4$ $61.3$ $10.4$ $64.5$
UMD 10 $2.8$ $28.7$ $2.2$ $52.1$ $3.5$ $31.9$ $8.8$ $63.7$ $8.3$ $67.7$
20 $2.4$ $30.1$ $1.8$ $66.0$ $3.5$ $33.4$ $5.0$ $74.0$ $7.0$ $84.4$
0 $4.1$ $34.1$ $4.4$ $54.1$ $5.0$ $36.6$ $14.3$ $74.5$ $11.5$ $66.1$
Unigram 10 $3.0$ $31.7$ $2.8$ $52.5$ $4.0$ $34.3$ $11.8$ $73.6$ $9.8$ $70.2$
20 $2.4$ $31.5$ $2.0$ $56.4$ $3.4$ $34.0$ $6.6$ $79.1$ $5.8$ $81.4$
In Table 1, we show the increase in perplexity on the training samples when the model is watermarked relative to the unwatermarked model. We observe that for Llama-30B, Unigram-Watermark induces a relative increase of $4.1$ in the minimum and $34.1$ in the average perplexity. Note that a relative increase of $4.1$ in perplexity for a sample makes it more than $4.3\times 10^{22}$ times less likely to be generated. This is based on a sample with only $32$ tokens, which is likely a lower bound since the number of tokens is typically larger than the number of words. We observe consistent results over several models and prompt lengths. For all experiments, unless otherwise specified, we use a fixed strength parameter $\delta=10$ for watermark methods and a fixed percentage of $50\%$green tokens. All the results are averaged over $5$ runs with different seeds for the watermark methods. We include additional results on WikiMIA-64, WikiMIA-128 and WikiMIA-256 in Tables 6, 7 and 8, respectively, in Appendix A.1. We observe that our findings are consistent across models and splits of WikiMIA.
In Figure 2 and Figure 4 from Appendix A.1, we study the influence of the strength of the watermark $\delta$ on the relative increase in both the minimum and average perplexity on the WikiMIA-32 training samples. In this experiment, we also consider a baseline of generating text freely to study the impact of watermarks on the quality of text relative to the impact on training samples’ generation (here, perplexity is computed by an unwatermarked model). All the results are averaged over $5$ runs with different seeds for the watermark methods. In the case of free generation, we generate $100$ samples for $5$ different watermarking seeds and average the results. The length of the generated samples is up to $42$ tokens, which is approximately $32$ words in the benchmark (on a token-to-word ratio of $4:3$). The results show an exponential increase in the perplexity of the training samples with the increase in watermark strength, while the generation quality is affected at a slower rate. This suggests that even if there is a trade-off between protecting the generation of text memorized verbatim and generating high-quality text, finding a suitable watermark strength for each particular application is possible.
@@ 4.1 Empirical Evaluation
0 $1.5$ $33.7$ $2.4$ $41.2$
UMD 10 $1.5$ $33.6$ $2.3$ $41.0$
@@
Takeaways. Watermarking significantly increases the perplexity of generating training texts, reducing verbatim memorization likelihood. This is achieved with only a moderate impact on the overall quality of generated text. This suggests that watermark strength can be effectively tailored to balance verbatim memorization and text quality for specific applications. Finally, we believe that our findings directly extend to the generation of copyrighted text verbatim, as this constitutes a form of verbatim memorization of the training data. Since copyrighted texts are not expected to be distributed significantly differently from the rest of the training data, the probability of generating copyrighted materials under watermarking is also likely to decrease. To confirm, we run similar experiments on a dataset containing copyrighted data (BookMIA) and include the results in the Table 2. Additionally, we consider finetuning Llama-7B [37] on BookMIA while controlling memorization by duplicating training samples. Detailed information about this experiment is provided in Appendix A.3. Finally, we extend our results to approximate memorization, which we define and discuss in Appendix A.3.
4.2 Theoretical analysis
Theorem 1.
For an LLM watermarked using a “hard” UMD scheme with a percentage of $\gamma$ green tokens, if $m\cdot\gamma^{n}<1$, then the probability of generating at least one copyrighted text from $T$ trials is lower than $m\cdot T\cdot\gamma^{n}$.
Proof. Given one sample $s=t_{1}\oplus t_{2}\oplus...\oplus t_{n}\in C_{D}$. For a “hard” watermarking scheme, the probability $P(s)$ of generating $s$ is smaller than the probability of each token $t_{i}$ to be on a green list. So, $P(s)<\gamma^{n}$. The probability of not generating any $s_{j}\in C_{D}$ is $P(\neg s_{1}\wedge\neg s_{2}\wedge...\wedge\neg s_{m})=\prod_{i=\overline{1,m}}(1-P(s_{i}))>(1-\gamma^{n})^{m}>1-m\gamma^{n}.$ Note that we used Bernoulli’s inequality at the end. The probability of not generating any $s_{j}\in C_{D}$ from $T$ trials is $>(1-m\gamma^{n})^{T}$, so it is larger than $1-m\cdot T\cdot\gamma^{n}$, again, by applying Bernoulli’s inequality. So, the probability of generating at least one copyrighted text from $T$ trials is lower than $1-(1-m\cdot T\cdot\gamma^{n})$ and hence lower than $m\cdot T\cdot\gamma^{n}$.
Example. Let’s consider a “hard” UMD watermarking scheme with $\gamma=0.5$. Let’s assume each copyrighted text is $100$ tokens, the model was trained on a dataset containing $10^{9}$ copyrighted texts and we make $10^{9}$ trials to generate copyrighted texts using the LLM. The probability to generate at least one copyrighted text out of these $10^{9}$ trials is $<10^{9}\cdot 10^{9}\cdot 0.5^{100}=\frac{10^{18}}{2^{100}}=\frac{1000^{6}}{1024^{10}}<\frac{1000^{6}}{1000^{10}}=1000^{-4}=10^{-12}$ and hence very low.
Theorem 2.
Proof. We include the proof in Appendix A.4.
Observation. Since the probability is reduced by at least $(1+\frac{2\epsilon}{2\epsilon+1})^{n}$ times in Theorem 2 then the probability of generating $s$ is lower than $(\frac{2\epsilon+1}{4\epsilon+1})^{n}$ (as the maximum probability of generating with the unwatermarked model is $1$). Hence, with the notations from Theorem 1, the probability of generating at least one copyrighted text from $T$ trials is lower than $m\cdot T\cdot(\frac{2\epsilon+1}{4\epsilon+1})^{n}$
Takeaways. Our theoretical analysis demonstrates that watermarking significantly reduces the probability of generating copyrighted text verbatim. For both a “hard” and “soft” UMD scheme, the upper bound for the likelihood of producing copyrighted content from a set of trials decreases exponentially with the length of the copyrighted texts.
5 Impact of Watermarking on Pretraining Data Detection
@@
Datasets. We revisit the WikiMIA benchmark as discussed in Section 4. We consider the full datasets, rather than the subset of samples that were part of the training for models we study. Additionally, we consider the BookMIA benchmark, which contains copyrighted texts.
@@
Metrics. We follow the prior work [31, 13] and report the AUC and AUC drop to study the detection performance of the MIAs. Note that this metric has the advantage of not having to tune the threshold for the detection classifier.
@@
Models. We conduct experiments on the same 5 LLMs as in Section 4. Additionally, for the Smaller Ref method that requires a smaller reference model along with the target LLM, we consider Llama-7B, Neo-125M, Pythia-70M, and OPT-350M as references.
5.1 Empirical evaluation

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in generating diverse and contextually rich text. However, concerns regarding copyright infringement arise as LLMs may inadvertently produce copyrighted material. In this paper, we first investigate the effectiveness of watermarking LLMs as a deterrent against the generation of copyrighted texts. Through theoretical analysis and empirical evaluation, we demonstrate that incorporating watermarks into LLMs significantly reduces the likelihood of generating copyrighted content, thereby addressing a critical concern in the deployment of LLMs. However, we also find that watermarking can have unintended consequences on Membership Inference Attacks (MIAs), which aim to discern whether a sample was part of the pretraining dataset and may be used to detect copyright violations. Surprisingly, we find that watermarking adversely affects the success rate of MIAs, complicating the task of detecting copyrighted text in the pretraining dataset. These results reveal the complex interplay between different regulatory measures, which may impact each other in unforeseen ways. Finally, we propose an adaptive technique to improve the success rate of a recent MIA under watermarking. Our findings underscore the importance of developing adaptive methods to study critical problems in LLMs with potential legal implications.
Introduction
Figure 1: Illustration of the effect of LLM watermarking on generation of copyrighted content. We observe that a moderately strong watermark ($\delta$ = 10) can make it more than $10^{20}$ times less likely for Llama-30B to generate copyrighted content.
@@ Introduction
Current lawsuits against AI companies for unauthorized use of copyrighted content (e.g., Andersen v. Stability AI Ltd, NYT v. OpenAI) highlight the urgent need for methods to address these challenges. In this paper, we focus on studying the effects of watermarking LLMs on two critical issues: (1) preventing the generation of copyrighted content, and (2) detecting copyrighted content in training data. We show that watermarking can significantly impact both the generation of copyrighted text and the detection of copyrighted content in training data.
Firstly, we observe that current LLM output watermarking techniques can significantly reduce the probability of LLMs generating copyrighted content, by tens of orders of magnitude. Our empirical results focus on two recent watermarking methods: UMD (Kirchenbauer et al. 2023) and Unigram-Watermark (Zhao et al. 2023). Both methods split the vocabulary into two sets (green and red) and bias the model towards selecting tokens from the green set by altering the logits distribution, thereby embedding a detectable signal. We provide both empirical and theoretical results to support our findings.
@@
Finally, we propose an adaptive method designed to enhance the success rate of a recent MIA (Shi et al. 2023) in detecting copyright violations under watermarking. This method applies a correction to the model’s output to account for the perturbations introduced by watermarks. By incorporating knowledge about the watermarking scheme, we improve the detection performance for pretraining data, counteracting the obfuscation caused by watermarking. Our contribution underscores the importance of continuously developing adaptive attack methodologies to keep pace with advances in defense mechanisms.
@@
The rest of the paper is organized as follows. In the “Related Work” section, we review prior research on LLM watermarking and copyright. The “Setup and Notations” section formally introduces the problems we study. We then present our first two contributions and introduce the adaptive version of the Min-KProb membership inference attack in the following three sections. Finally, we provide concluding remarks in the last section. Additional experiments, theoretical results, and a discussion on the limitations of our work are included in the appendix
Related Work
@@
Watermarks for LLMs. Language model watermarking techniques embed identifiable markers into output text to detect AI-generated content. Recent strategies incorporate watermarks during the decoding phase of language models (Zhao et al. 2023; Kirchenbauer et al. 2023). Aaronson (2023) develops the Gumbel watermark, which employs traceable pseudo-random sampling for generating subsequent tokens. Kirchenbauer et al. (2023) splits the vocabulary into red and green lists according to preceding tokens, biasing the generation towards green tokens. Zhao et al. (2023) employs a fixed grouping strategy to develop a robust watermark with theoretical guarantees. Liu et al. (2024) proposes to generate watermark logits based on the preceding tokens’ semantics rather than their token IDs to boost the robustness. Kuditipudi et al. (2023) and Christ, Gunn, and Zamir (2023) explore watermark methods that do not change the output textual distribution.
@@
Copyright. Copyright protection in the age of AI has gained importance, as discussed by Ren et al. (2024). Vyas, Kakade, and Barak (2023) addresses content protection through near access-freeness (NAF) and developed learning algorithms for generative models to ensure compliance under NAF conditions. Prior works focus on training algorithms to prevent copyrighted text generation (Vyas, Kakade, and Barak 2023; Chu, Song, and Yang 2024), whereas our work emphasizes lightweight, inference-time algorithms. Other works have studied copyright in machine learning from a legal perspective. Hacohen et al. (2024) utilizes a generative model to determine the generic characteristics of works to aid in defining the scope of copyright. Elkin-Koren et al. (2023) demonstrates that copying does not necessarily constitute copyright infringement and argues that existing detection methods may detract from the foundational purposes of copyright law.
Additionally, we include a discussion on memorization and membership inference in the appendix.
Setup and Notations
Definitions
@@
Verbatim Memorization of Copyrighted Content. For a fixed $k\in\mathbb{N}$, Carlini et al. (2022b) defines a string $s$ as being memorized by a model if $s$ is extractable with a prompt $p$ of length $k$ using greedy decoding and the concatenation $p\oplus s\in D$. We adopt a similar definition for verbatim memorization of copyrighted content but employ a continuous metric to measure it. Specifically, we measure verbatim memorization of a text $c\in C$ using the perplexity of the model on the copyrighted text $c_{p}$ when given the prefix $p$ as a prompt (where $c_{p}$ represents the text $c$ after removing its prefix $p$). Note that for $c_{p}=c_{p}^{(1)}\oplus c_{p}^{(2)}\oplus\dots\oplus c_{p}^{(n)}$ we compute the perplexity using the following formula $\texttt{perplexity}(c_{p}|p)=\big{(}\prod_{i=1}^{n}\mathbb{P}(c_{p}^{(i)}|p\oplus c_{p}^{(0)}\oplus c_{p}^{(1)}\oplus\dots\oplus c_{p}^{(i-1)}\big{)}^{-\frac{1}{n}}$, where $c_{p}^{(0)}$ is the empty string. In our experiments, $p$ is either an empty string or the first $10$, $20$, or $100$ tokens of $c$. Lower perplexity thereby indicate higher levels of memorization.
MIAs for Copyrighted Training Data Detection. MIAs are privacy attacks aiming to detect whether a sample was part of the training set. We define an MIA for copyrighted data as a binary classifier $A(\cdot)$, which ideally outputs $A(x)=1,\forall x\in C_{D}$ and $A(x)=0,\forall x\in C-C_{D}$. In practice, $A(\cdot)$ is defined by thresholding a metric (e.g., perplexity), i.e., $A(x)=1,\forall x$ such that perplexity$(x)<t$ and 00, otherwise. Since the threshold $t$ needs to be set, prior work (Shi et al. 2023) uses AUC (Area Under the ROC Curve) as an evaluation metric which is independent of $t$. Note that we employ the same metric in our experiments.
LLM Watermarking. Watermarking LLMs consists of introducing signals during its training or inference that are difficult to detect by humans without the knowledge of a watermark key but can be detected using an algorithm if the key is known. We focus our paper on recent methods that employ logits distribution changes as a way of inserting watermark signals during the decoding process (Kirchenbauer et al. 2023; Zhao et al. 2023).
MIAs
@@
Smaller Ref, Lowercase and Zlib (Carlini et al. 2021). Smaller Ref is defined as the ratio of the log-perplexity of the target LLM on a sample to the log-perplexity of a smaller reference LLM on the same sample. Lowercase represents the ratio of the log-perplexity of the target LLM on the original sample to the log-perplexity of the LLM on the lowercase version of the sample. Zlib is defined as the ratio of the log-perplexity of the target LLM on a sample to the zlib entropy of the same sample.
@@
Min-KProb (Shi et al. 2023). This heuristic computes the average of the minimum $K\%$ token probabilities outputted by the LLM on the sample. Note that this method requires tuning $K$, so in all our experiments we chose the best result over $K\%\in\{5\%,10\%,20\%,30\%,40\%,50\%,60\%\}$.
LLM Watermarking Methods
@@
UMD (Kirchenbauer et al. 2023) splits the vocabulary into two sets (green and red) and biases the model towards the green tokens by altering the logit distribution. The hash of the previous token’s ID serves as a seed for a pseudo-random number generator used to split the vocabulary into these two groups. For a “hard” watermark, the model is forced not to sample from the red list at all. For a “soft” watermark, a positive bias $\delta$ is added to the logits of the green tokens before sampling. We focus our empirical evaluation on “soft” watermarks as they are more suitable for LLM deployment due to their smaller impact on the quality of the generated text.
@@
Unigram-Watermark (Zhao et al. 2023) employs a similar approach of splitting the vocabulary into two sets and biasing the model towards one of the two sets. However, the split remains consistent throughout the generation. This choice is made to provide a provable improvement against paraphrasing attacks (Krishna et al. 2024).
Watermarking LLMs Prevents
Copyrighted Text Generation
@@
In this section, we study the effect of LLM watermarking techniques on verbatim memorization. We discuss the their implications for preventing copyrighted text generation.
@@
Datasets. We consider $4$ versions of the WikiMIA benchmark (Shi et al. 2023) with 32, 64, 128, and 256 words in each sample and only consider the samples that were very likely part of the training set of all the models we consider (labeled as $1$ in Shi et al. (2023)). We consider these subsets as a proxy for text that was used in the training set, and the model may be prone to verbatim memorization. From now on, we refer to this subset as the “training samples” or “training texts”. Similarly, we consider BookMIA dataset (Shi et al. 2023), which contains samples from copyrighted books.
@@
Metric. We measure the relative increase in perplexity on the generation of training samples by the watermarked model compared to the original model. We report the increase in both the minimum and average perplexity over the training samples. Note that a large increase in perplexity corresponds to a large decrease in the probability of generating that specific sample, as shown later in this section. When computing the perplexity, we prompt the model with an empty string, the first 10, and the first 20 tokens of the targeted training sample, respectively. In the BookMIA dataset, we designate the initial 100 or 256 tokens as the prompt. This is because each BookMIA sample contains 512 words, which is larger than the sample size in WikiMIA.
@@
Models. We conduct our empirical evaluation on $5$ recent LLMs: Llama-30B (Touvron et al. 2023), GPT-NeoX-20B (Black et al. 2022), Llama-13B (Touvron et al. 2023), Pythia-2.8B (Biderman et al. 2023) and OPT-2.7B (Zhang et al. 2022).
Empirical Evaluation
@@
UMD 0 $3.3$ $31.2$ $4.9$ $34.3$
10 $2.8$ $28.7$ $3.5$ $31.9$
20 $2.4$ $30.1$ $3.5$ $33.4$
Unigram 0 $4.1$ $34.1$ $5.0$ $36.6$
10 $3.0$ $31.7$ $4.0$ $34.3$
20 $2.4$ $31.5$ $3.4$ $34.0$
In Table 1, we show the increase in perplexity on the training samples when the model is watermarked relative to the unwatermarked model. We observe that for Llama-30B, Unigram-Watermark induces a relative increase of $4.1$ in the minimum and $34.1$ in the average perplexity. Note that a relative increase of $4.1$ in perplexity for a sample makes it more than $4.3\times 10^{22}$ times less likely to be generated. This is based on a sample with only $32$ tokens, which is likely a lower bound since the number of tokens is typically larger than the number of words. We observe consistent results over several models and prompt lengths. For all experiments, unless otherwise specified, we use a fixed strength parameter $\delta=10$ for watermark methods and a fixed percentage of $50\%$ green tokens. All the results are averaged over $5$ runs with different seeds for the watermark methods. We include additional results on WikiMIA-64, WikiMIA-128 and WikiMIA-256 in Tables 7, 8 and 9, respectively, in the appendix. We observe that our findings are consistent across models and splits of WikiMIA. Finally, we include the complete version of Table 1 in the appendix (Table 6), which shows results for additional models and random logit perturbations with the same strength as the watermarking methods. Overall, the additional results are consistent with our previous findings.
Llama-30B Llama-13B
P. Min. Avg. Min. Avg.
UMD 0 $1.5$ $33.7$ $2.4$ $41.2$
10 $1.5$ $33.6$ $2.3$ $41.0$
@@
In Figure 2, as well as Figure 5 from the appendix, we study the influence of the strength of the watermark $\delta$ on the relative increase in both the minimum and average perplexity on the WikiMIA-32 training samples. In this experiment, we also consider a baseline of generating text freely to study the impact of watermarks on the quality of text relative to the impact on training samples’ generation (here, perplexity is computed by an unwatermarked model). All the results are averaged over $5$ runs with different seeds for the watermark methods. In the case of free generation, we generate $100$ samples for $5$ different watermarking seeds and average the results. The length of the generated samples is up to $42$ tokens, which is approximately $32$ words in the benchmark (on a token-to-word ratio of $4:3$). The results show an exponential increase in the perplexity of the training samples with the increase in watermark strength, while the generation quality is affected at a slower rate. This suggests that even if there is a trade-off between protecting the generation of text memorized verbatim and generating high-quality text, finding a suitable watermark strength for each particular application is possible. Examples of generated samples at varying watermark strengths are provided in the appendix.
Approximate Memorization. Informally, we consider a training sample approximately memorized by a model if, given its prefix, it is possible to generate a completion that is similar enough to the ground truth completion. In our experiments, we use models fine-tuned on a subset of BookMIA (details provided in the appendix) and we consider Normalized Edit Similarity (referred to as edit similarity from now on) and BLEU score as similarity measures, as in (Ippolito et al. 2023). Note that we consider both word-level and token-level variants for the BLEU score. The range for each metric is between 0 and 1, where values close to 1 represent similar texts. In all experiments, since all the samples are 512 words long, we consider the first 256 words as the prefix and the last 256 words as the ground truth completion. We present the results for edit similarity with the UMD watermark in Figure 3, and the complete results—using all metrics and including the Unigram watermark—in Figure 7, averaged over 20 runs with different random seeds. Note that the duplication factor (shown on x-axis) represents the number of times the target copyrighted text is duplicated. We observe that for high levels of memorization, a strong watermark significantly reduces the similarity between the generated completion and the ground truth (copyrighted) one.
Figure 3: Edit similarity between the generated completion and the ground truth when considering different watermark strengths and memorization levels.
@@
Takeaways. Watermarking significantly increases the perplexity of generating training texts, reducing verbatim memorization likelihood. This is achieved with only a moderate impact on the overall quality of generated text. This suggests that watermark strength can be effectively tailored to balance verbatim memorization and text quality for specific applications. Finally, we believe that our findings on WikiMIA—which does not necessarily contain copyrighted data—directly extend to the generation of copyrighted text verbatim, as this constitutes a form of verbatim memorization of the training data. To confirm, we run similar experiments on a dataset containing copyrighted data (BookMIA) and include the results in the Table 2. Additionally, we consider finetuning Llama-7B (Touvron et al. 2023) on BookMIA while controlling memorization by duplicating training samples. Detailed information about this experiment is provided in the appendix.
Impact of Watermarking on
Pretraining Data Detection
@@
Datasets. We revisit the WikiMIA benchmark as discussed in the previous section. We consider the full datasets, rather than the subset of samples that were part of the training for models we study. Additionally, we consider the BookMIA benchmark, which contains copyrighted texts.
@@
Metrics. We follow the prior work (Shi et al. 2023; Duarte et al. 2024) and report the AUC and AUC drop to study the detection performance of the MIAs. Note that this metric has the advantage of not having to tune the threshold for the detection classifier.
@@
Models. We conduct experiments on the same LLMs as in the previous section. Additionally, for the Smaller Ref method that requires a smaller reference model along with the target LLM, we consider Llama-7B, Neo-125M, Pythia-70M, and OPT-350M as references.
Empirical Evaluation